OpenStudy (anonymous):

What Improper fraction falls in between 6 1/4 and 6 1/3

sammixboo (sammixboo):

Take the fractions \(\sf \large \color{magenta}{6\frac{1}{4}}\) and \(\sf \large \color{magenta}{6\frac{1}{3}}\) and make them have a common denominator of 24. \(\sf \large \color{magenta}{6\frac{1 x 6}{4 x 6}}\) ---- \(\sf \large \color{magenta}{6\frac{6}{24}}\) \(\sf \large \color{magenta}{6\frac{1 x 8}{3 x 8}}\) ---- \(\sf \large \color{magenta}{6\frac{8}{24}}\) What fraction is between \(\sf \large \color{magenta}{6\frac{6}{24}}\) and \(\sf \large \color{magenta}{6\frac{8}{24}}\)

sammixboo (sammixboo):

\(\sf \large \color{magenta}{6\frac{7}{24}}\) is in between those fractions. Do you know what \(\sf \large \color{magenta}{6\frac{7}{24}}\) in improper form is?

OpenStudy (nincompoop):

sammi, you're quite suspicious

sammixboo (sammixboo):

Am I? o_o

sammixboo (sammixboo):

<_<..... >_>.....

OpenStudy (nincompoop):

first: convert them to improper fractions

sammixboo (sammixboo):